In 2014, a major drug company agreed to give a not-for-profit private college $1,700,000 to perform...

  1. In 2014, a major drug company agreed to give a not-for-profit private college $1,700,000 to perform testing of a new drug. An advance payment of $700,000 was received in 2014. The college was to receive $4,000 per individual test. In 2014, the college completed 100 tests. How much revenue should the college report for 2014?

            A) $     - 0 - .

            B)   $   400,000.

            C)   $   700,000.

            D) $1,700,000.

Answer: B please show the calculations

2.  A private foundation made a multi-year pledge to a private college on December 31, 2013, the last day of the fiscal year. The pledge was to pay $12,000 per year each year for five years, beginning on December 31, 2014. The discount rate is 6%. The present value of five payments of $12,000 is $50,548. The present value of four payments of $12,000 is $41,581. No purpose or plant restrictions were involved.

1. The private college would:

            A) Record interest revenue of $3,033 in 2013.

            B)   Record interest revenue of $3,033 in 2014.

            C)   Record contribution revenue of $3,033 in 2013.

D) Record contribution revenue of $3,033 in 2014.

Answer: D Show the calculation
Answer #1

1) Since only 100 tests are performed in 2014, only revenue for these 100 tests will be recognised in 2014.

Revenue per test = $4,000

therefore, Revenue for 100 tests = 100 *4000 = $400,000. Answer

2) Contribution Revenue = 50548*0.06 = $3,033